I am stuck on these types of problems. Please assist.
Solve by graphing. Indicate wether each system is independent, inconsistent, or dependent.
3x-y=4
3x-y=0

Answer by tutorcecilia(2152)

3x-y=4
3x-y=0
_______ subtract the equations
0+0=4 [this is an untrue statement]
.
The system is inconsistent and independent because there are no points in common between the two lines. Also, the lines are parallel since they have the same slope.